# Wall-Proximity Matters: Understanding the Effect of Device Placement with Respect to the Wall for Indoor Wi-Fi Sensing

### v2

@@ VI-B Evaluating Sensing Coverage of the Proposed Model through Stationary Crowd Counting
To validate the sensing coverage aspects of the proposed model, we conducted experiments tailored specifically for stationary crowd counting applications. The experimental setup in Sections VI-B and VI-C included five participants of varying heights and weights, consisting of one female and four males, aged 24 to 26 years. Each experiment consisted of three trials, with participants positioned at a total of 15 distinct locations distributed randomly throughout the entire space. In these experiments, we varied the wall-device distance while keeping other factors constant, such as the participants’ positions and the total number of participants. We validated the model and the simulated sensing boundary by correlating them with the counting accuracy. Higher accuracy indicates better performance, implying a larger sensing coverage area in the deployment. While the real sensing area expands beyond the simulated area due to advanced signal preprocessing methods, the trends observed with varying device positions should still align with the simulated model.
@@
In this experiment, we aim to verify the impact of the wall-device distance on the sensing coverage, as discussed in Section V-A. The device deployment settings are the same as in Section 10, as depicted in Fig. 11. Fig. 14 presents the accuracy results (line charts) obtained for the three cases, along with the corresponding simulation-based sensing areas (bar chart). We observe that the experimental results across all three models align well with the simulation trend. Specifically, as the distance between the wall and the device decreases within a limited range, the indoor sensing area increases. This finding confirms the relationship between the wall-device distance and the extent of the sensing coverage, as predicted by the simulation analysis.
Figure 14: Evaluation results of the effect of the wall-device distance without interference.
@@
As mentioned in Section IV-B, when the device is close to the wall, the sensing area on the other side of the wall also increases, which will have a negative effect on indoor sensing applications. This experiment is designed to verify this property. The experiment setup is the same as Experiment 1 except that there are people outside the room and close to the wall. In Fig. 15, the green bars represent the sensing areas beyond the wall and the three lines show the accuracy results across all three models for people counting within the room. As the Tx moves closer to the wall, the sensing area on the other side of the wall extends, which indicates a large interference area in this experiment. As can be seen from Fig. 15, with an increase in the wall-device distance, the accuracy increases. This trend verifies the negative effect of a short wall-device distance when there are interferers beyond the wall.
Figure 15: Evaluation results of the effect of the wall-device distance with interference.
@@
To evaluate the effect of the transmitter-receiver distance on the proposed model, we conducted experiments in a meeting room with a fixed distance of 1.2 m between the wall and the Tx. As depicted in Fig. 16, we varied the distance between the transceivers from 0 m to 6 m at a step size of 1 m. The results, shown in Fig. 17, indicate that the accuracy initially increases and then slightly decreases with the increased distance between the Tx and Rx. This trend is consistent across all the three models, and it aligns with the simulated sensing area presented in the bar chart. Notably, an optimal distance of 3 m emerged, resulting in the largest sensing area and the highest accuracy for stationary crowd counting in this particular environment. These experimental results serve to validate the property of the proposed model detailed in Section 6.
While the phenomena observed in Experiment 3 could also be explained by the existing model [17], it fails to accurately represent the outputs from Experiments 1 and 2, as the previous model does not consider the influence of walls on dynamic sensing capabilities. In contrast, our model uniquely demonstrates the effects of wall-device distance, offering a more comprehensive understanding of deployment strategies in real-world scenarios.
Figure 17: Evaluation results of the effect of the Tx-Rx distance.
@@ VI-C Boosting the Sensing Coverage of Stationary Crowd Counting
    In the second scenario, we maintain the same settings as in the first scenario but in a different environment—a classroom with a different size compared to the meeting room. As a result, the distance between the transceivers is distinct from the first scenario. This scenario is illustrated in Fig. 21.
@@ VII Discussion
Multiple walls. The proposed model currently considers the effect of only a single wall on dynamic power, whereas in reality, indoor spaces typically have four walls. To enhance the accuracy of the model, it would be beneficial to incorporate the superposition of the impacts of all surrounding walls.
@@
Other factors for optimal device deployment. In this study on stationary crowd counting, targets were randomly positioned throughout the space, without considering the impact of their locations on sensing capabilities. It is important to note that in real-world scenarios, optimal device performance may not always involve placing devices near walls, as other factors, such as target positioning and obstacles, can significantly influence sensing outcomes. Future work could involve conducting experiments in a wider range of environments, incorporating various room sizes, layouts, and wall materials. By integrating the positions of targets and other relevant factors into deployment strategies, we can improve our understanding of how device placement affects sensing performance. This approach ultimately leads to more accurate and practical applications.
@@ VIII Conclusion
Our work highlights the significant impact of both the wall-device distance and transmitter-receiver distance on Wi-Fi sensing coverage in indoor environments. By introducing a theoretical model, we have successfully demonstrated the relationship between wall-device distances and sensing coverage. The robustness of our proposed model was assessed through the execution of experiments focusing on respiration monitoring and stationary crowd counting across diverse environmental settings. The insights gained from this model enable us to strategically plan device placement, ultimately expanding the sensing coverage in practical deployments of wall-reflection sensing systems. This research successfully bridges the gap between theoretical understanding and practical implementation, paving the way for improved real-world deployment of such systems. Our findings contribute to advancing the field of Wi-Fi sensing and offer valuable guidance for optimizing sensing performance in indoor environments.

### v4

@@ VI-B Evaluating Sensing Coverage of the Proposed Model through Stationary Crowd Counting
To validate the sensing coverage aspects of the proposed model, we conducted experiments tailored specifically for stationary crowd counting applications. The experimental setup in Sections VI-B and VI-C included five participants of varying heights and weights, consisting of one female and four males, aged 24 to 26 years. Each experiment consisted of three trials, with participants positioned at a total of 15 distinct locations distributed randomly throughout the entire space. In these experiments, we varied the wall-device distance while keeping other factors constant, such as the participants’ positions and the total number of participants. Since crowd counting involves multiple static targets simultaneously, the SSNR metric is not applicable in this scenario. Instead, we validate the model and its simulated sensing coverage by correlating the estimated sensing areas with the counting accuracy. It should be noted that the simulated sensing area derived from the SSNR-based model serves as a theoretical guideline rather than a strict predictor of recognition accuracy. Higher accuracy indicates better performance and implies a larger effective sensing coverage under ideal conditions. However, this relationship may not always hold in practice, as larger coverage could also introduce greater vulnerability to external interference. This trade-off will be further examined in subsequent experiments. While the real sensing area may extend beyond the simulated boundary due to advanced signal preprocessing, the overall trends observed across varying device positions should still align with the model predictions.
@@
In this experiment, we aim to verify the impact of the wall-device distance on the sensing coverage, as discussed in Section V-A. The device deployment settings are the same as in Section 10, as depicted in Fig. 11. Fig. 13a presents the accuracy results (line charts) obtained for the three cases, along with the corresponding simulation-based sensing areas (bar chart). We observe that the experimental results across all three models align well with the simulation trend. Specifically, as the distance between the wall and the device decreases within a limited range, the indoor sensing area increases. This finding confirms the relationship between the wall-device distance and the extent of the sensing coverage, consistent with the simulated property referred to as Factor I in Section V-A.
Figure 13: Evaluation results of the effect of the wall-device distance on sensing performance. (a) Without interference. (b) With interference. Bars: sensing area; lines: accuracy.
@@
As mentioned in Section IV-B, when the device is close to the wall, the sensing area on the other side of the wall also increases, which will have a negative effect on indoor sensing applications. This experiment is designed to verify this property. The experiment setup is the same as Experiment 1 except that additional people are positioned outside the room, adjacent to the wall, introducing external interference. In Fig. 13b, the green bars represent the sensing areas beyond the wall and the three lines show the accuracy results across all three models for people counting within the room. As the Tx moves closer to the wall, the sensing area on the other side of the wall extends, which indicates a large interference area in this experiment. As can be seen from Fig. 13b, with a decrease in the wall-device distance, the accuracy also decreases. This trend verifies the negative effect of a short wall-device distance when there are interferers beyond the wall.
@@
To evaluate the effect of the transmitter-receiver distance on the proposed model, we conducted experiments in a meeting room with a fixed distance of 1.2 m between the wall and the Tx. As depicted in Fig. 14, we varied the distance between the transceivers from 0 m to 6 m at a step size of 1 m. The results, shown in Fig. 15, indicate that the accuracy initially increases and then slightly decreases with the increased distance between the Tx and Rx. This trend is consistent across all the three models, and it aligns with the simulated sensing area presented in the bar chart. Notably, an optimal distance of 3 m emerged, resulting in the largest sensing area and the highest accuracy for stationary crowd counting in this particular environment. These experimental results serve to validate the property of the proposed model referred to as Factor II in Section 6.
While the phenomena observed in Experiment 3 could also be explained by the existing model [13], it fails to capture the key observations from Experiments 1 and 2, as the previous model does not consider the influence of walls on dynamic sensing capabilities. In contrast, our model uniquely demonstrates the effects of wall-device distance, offering a more comprehensive understanding of deployment strategies in real-world scenarios. This deeper understanding suggests that while closer proximity to the wall can improve sensing coverage, it may be counterproductive in environments with potential interference sources. Therefore, deployment decisions should balance sensing range expansion with environmental robustness, particularly in complex indoor settings.
Figure 15: Evaluation results of the effect of the Tx-Rx distance. Bars: sensing area; lines: accuracy.
@@ VI-C Boosting the Sensing Coverage of Stationary Crowd Counting
    In the second scenario, we maintain the same settings as in the first scenario but in a different environment, a classroom with a different size compared to the meeting room. As a result, the distance between the transceivers is distinct from the first scenario. This scenario is illustrated in Fig. 19.
@@
Overall, the experimental observations in Sections VI-A and VI-B exhibit strong consistency with the simulation-based analyses presented in Section V. As illustrated in Figs. 12, 13, and 15, the simulated effects of wall–Tx distance and Tx–Rx distance on SSNR and sensing area align well with the observed trends in the corresponding real-world performance metrics, including MAE of respiratory rate and crowd counting accuracy. These results collectively validate the theoretical models and simulation assumptions adopted in this work. Furthermore, the extended deployment scenarios evaluated in Section VI-C demonstrate that these modeled properties can be effectively applied in more complex real-world settings, further supporting their practical value in boosting sensing coverage.
@@ VII Discussion
Multiple environments. Although the experiments were conducted in three representative indoor environments (a meeting room and two classrooms) with five participants, these setups were carefully designed to ensure controlled and repeatable measurements across different transceiver placements. The proposed framework is general in formulation and can be extended to other room geometries and materials; however, further validation in more diverse environments (e.g., furniture-rich or metallic spaces) will be pursued in future work to strengthen the generalization of the findings.
@@
Other factors for optimal device deployment. In this study on stationary crowd counting, targets were randomly positioned throughout the space, without explicitly considering the impact of their locations on sensing capabilities. It is important to note that in real-world scenarios, optimal device performance may not always involve placing devices near walls, as other factors, such as target positioning, obstacles, and device movement direction, can also influence sensing outcomes. Future work could involve conducting experiments in a wider range of environments, incorporating various room sizes, layouts, and wall materials, as well as directionally diverse device movements. By integrating these factors into deployment strategies, we can further generalize the proposed framework and deepen our understanding of how device placement affects sensing performance, thereby supporting more accurate and practical applications.
@@ VIII Conclusion
Our work highlights the significant impact of both the wall-device distance and transmitter-receiver distance on Wi-Fi sensing coverage in indoor environments. By introducing a theoretical model, we have successfully demonstrated the relationship between wall-device distances and sensing coverage. The performance of our proposed model was evaluated through experiments on respiration monitoring and stationary crowd counting across representative indoor settings. The insights gained from this model enable us to strategically plan device placement, ultimately expanding the sensing coverage in practical deployments of wall-reflection sensing systems. This research successfully bridges the gap between theoretical understanding and practical implementation, paving the way for deployment-aware Wi-Fi sensing in typical indoor scenarios. Our findings contribute to advancing the field of Wi-Fi sensing and offer valuable guidance for optimizing sensing performance in realistic indoor environments